An original rectangle has a length of 8 and a width of 24. What happens to the area if the new width is one half the original width? Verify your answer.
Answer by CharlesG2(834)

An original rectangle has a length of 8 and a width of 24. What happens to the area if the new width is one half the original width? Verify your answer.
original rectangle Area = Length * Width = 8 * 24 = 192
1/2 * 24 = 12
new rectangle Area = 8 * 12 = 96
the Area halves